What will be the resulting normality of NaCl solution formed on mixing 200ml of decimolar NaOH with 40ml of semimolar HCl?

Grade:11

1 Answers

Sujit Kumar
5 years ago
Decimolar solution means 0.1M
Semimolar solution means 0.5M
 
NaOH + HCl → NaCl + H2O
 
No. Of millimoles of NaOH = 200*0.1 = 20
No. Of millimoles of HCl = 40*0.5 = 20
 
→ No. Of millimoles of NaCl formed = 20
→ Normality = (no. Of millimoles * n-factor)/volume of sol.(in mL)
→ Normality = (20*1)/240                           (240mL is the total volume i.e 200+40, 1 is the n-factor of NaCl)
→ Normality = 0.0833 N
 
Ans: 0.0833 N